Calculate the maximum monochromatic emissive power at 288K?
1
Expert's answer
2021-09-02T10:41:37-0400

According to Wien's displacement  law, we have

"\lambda=\frac bT."


The maximum monochromatic emissive power is


"E=\frac{2hc^2}{\lambda^5}·\frac{1}{e^{\frac{hc}{\lambda kT}}-1}."

Substitute the wavelength:


"E=\frac{2hc^2T^5}{b^5}·\frac{1}{e^{\frac{hc}{b k }}-1}=8.12·10^6\text{ W/m}^2."